# On the Reasoning Abilities of Masked Diffusion Language Models

### v1

@@ Abstract
Masked diffusion models (MDMs) for text offer a compelling alternative to traditional autoregressive language models. Parallel generation makes them efficient, but their computational capabilities and the limitations inherent to their parallelism remain largely unexplored. To this end, we characterize what types of reasoning problems MDMs can provably solve and how efficiently. We do this by connecting MDMs to the well-understood reasoning frameworks of chain of thought (CoT) and padded looped transformers (PLTs) in the finite-precision log-width setting: We show that MDMs and polynomially-padded PLTs are, in fact, equivalent in this setting, and that MDMs can solve all problems that CoT-augmented transformers can. Moreover, we showcase classes of problems (including regular languages) for which MDMs are inherently more efficient than CoT transformers, where parallel generation allows for substantially faster reasoning.
@@ 1 Introduction
Many complex problems can be decomposed into smaller, independent sub-problems, making them naturally suited for parallel computation. For example, we can compute the value of a mathematical expression by evaluating its sub-expressions independently and combining the results (see Fig.˜2). However, dominant autoregressive language models (LMs) tackle these problems sequentially. Methods like chain of thought (CoT), for instance, generate solutions one step at a time, failing to capitalize on the underlying parallel structure. Parallel generation by masked diffusion models (MDMs) offers a compelling alternative. Recent advances have positioned MDMs as a viable contender to autoregressive LMs in language modeling, code generation, and even molecule design (Lou et al., 2024; Zhang et al., 2025; Sun et al., 2025). However, the fundamental reasoning capabilities of MDMs remain poorly understood, which limits the extent to which we can leverage their potential and apply them to appropriate tasks. This work bridges that gap by providing the first formal characterization of the expressivity of MDMs, clarifying their fundamental computational strengths and weaknesses.
@@
Figure 1: A summary of masked diffusion model expressivity. The colored nodes in the top row correspond to transformer-based computational models and the languages recognized by them. The gray nodes in the bottom row correspond to languages in (${\mathtt{L}}\textnormal{-uniform}$}) classical complexity classes. ${{{\mathcal{X}}}}\hookrightarrow{{{\mathcal{Y}}}}$ indicates the inclusion of $\mathcal{X}$}}} in $\mathcal{Y}$}}}, and ${{{\mathcal{X}}}}\leftrightarrow{{{\mathcal{Y}}}}$ indicates equality. Dashed arrows represent strict inclusions. Red arrows denote novel results. $\mathtt{Reg}$ refers to all regular languages.
@@
Our analysis builds upon the study of LM expressivity—the formal characterization of the problems whose solution the neural architecture of an LM (with appropriate parameters) can express. While this field comprehensively describes autoregressive LMs, its findings do not directly apply to MDMs due to their fundamentally different, non-sequential processing of text, leaving the theoretical studies of these two paradigms largely in isolation. Prior theoretical work on MDMs has focused on the limitations of their factorized backward process and its convergence properties. This research has shown that while MDMs can approximate $n$-gram LMs with constantly many denoising steps, the number of steps must grow linearly with the string length even for simple LMs such as probabilistic regular languages (Feng et al., 2025; Li & Cai, 2025). [^1] We provide a detailed overview of related work in App. A. However, the asymptotic nature of these findings, together with the strict assumptions on the theoretical model, makes it difficult to draw concrete conclusions about the practical reasoning capabilities of MDMs, leaving a critical gap: A theoretical framework for MDMs that is both (1) formally rigorous enough to provide a comprehensive picture of how MDMs can use the combination of parallelism and iterative refinement for formal reasoning, and (2) faithful to how they are implemented in practice. Our work introduces such a framework, providing a tight characterization of their reasoning capabilities.
@@
We refer to the fact that CoT cannot take advantage of this parallelism as the sequentiality bottleneck of CoT. Takeaway˜2 highlights the potential efficiency gains of parallel CoT that generates multiple symbols at once. Our analysis, in fact, identifies a tighter and more natural connection between MDMs and this variant, which we refer to as pCoT. The parallelism of MDMs also facilitates a close connection with looped and padded transformers, where looping naturally maps to denoising steps and padding tokens to the generated tokens of an MDM. We find the class of problems solvable by MDMs is, in fact, precisely equivalent to that solvable by padded looped transformers.
@@
We summarize these takeaways and their relationships in Fig.˜1. Our proofs emphasize the affordances and difficulties in solving reasoning problems with MDMs. We find, for example, that the discrete nature of text generation, which serves as a communication channel between individual denoising steps, limits the amount of information that can be passed between steps. This necessitates extra output space to store the intermediate computations and is analogous to the discrepancy between fixed-precision and log-precision transformers (Li et al., 2024), and the difference between the classes $\mathtt{AC}$}} and $\mathtt{TC}$}}, the class of circuits with threshold gates. We find that positional encodings that carry information not computable by transformers themselves are crucial in locating this information. We also observe that, while unmasked attention makes MDM attention patterns flexible, it complicates left-to-right processing, which is often natural in human language—we find that exact implementation of causal masking requires quadratically more output space.
@@ 2.2.1 Our idealization of masked diffusion models
We aim to understand the expressivity of the reverse process. To this end, we introduce an idealization that captures its key aspects—iterative unmasking and infilling—and provides a principled lens for understanding the expressivity of practical MDMs and comparing them to well-known paradigms such CoT. Here, we describe the high-level ideas; see §˜B.7 for the full formal model.
@@
We denote the class our idealized MDMs by $\mathtt{MDM}$}. For a more detailed discussion, including the theoretical connection of our idealization to existing MDM variants, see App.˜C.
@@ 3.2 MDMs and CoT can (inefficiently) simulate each other
The simulation incurs a quadratic blow-up in the padding length. This is not due to an inherent feature of the diffusion process but rather the challenge of simulating masked attention with unmasked one (cf. Lem.˜D.14). [^13] To the best of our knowledge, Lem. D.14 is the first result showing how to simulate masked attention with unmasked attention and might be of interest in its own right. In particular, if the MDM transformer is causally masked, the blow-up disappears. Moreover, if the unmasked transformer can simulate masking more efficiently (with, for example, more expressive scoring functions), the blow-up can be alleviated. [^14] This is, for example, used by Saunshi et al. (2025, Thm 5.4) to show that unmasked PLTs can simulate CoT with no blow-up in the padding using a masking function—an additional step in the computation of attention scores that allows for zeroing out of irrelevant keys—and with linearly-increasing width. While Lem.˜D.14 could possibly be improved, it is interesting to note that the seemingly more general unmasked nature of MDMs might negatively impact their ability to align with human-oriented sequential processing captured by causal masking, which could provide a useful inductive bias for the masked models. [^15] Interestingly, some existing work finds that MDMs that decode based on the most confident symbols tend to decode autoregressively (Gong et al., 2025). In this sense, the unmasked nature of MDMs could be seen as a hurdle that the model has to overcome to eventually rely on more autoregressive generation. This further motivates the development of hybrid models that combine autoregressive generation of entire blocks with non-autoregressive infilling within the blocks (Nie et al., 2025; Arriola et al., 2025; Song et al., 2025, inter alia).
@@ 4 Discussion
In other words, these problems, in general, require a “CoT-style” step-by-step sequential solution. Due to the overhead introduced by unmasked attention of MDMs (such as the inability to store KV-cache), such problems are more efficiently solved by standard autoregressive CoT transformers.
@@
We strive to compare different paradigms fairly by analyzing a specific implementation of MDMs—one based on a specific idealization of transformers. Some inherent differences between the frameworks, however, remain. One is the dichotomy between using causal masking for CoT and unmasked transformers for MDMs and PLTs. We focus on unmasked models to analyze the more natural and popular implementations rather than artificially constraining MDMs and PLTs to causal masking. Another impactful decision is the nature of positional encodings (PEs). While we assume relatively simple PEs standard in theoretical literature (in particular, logspace-computable, cf. §˜B.4), we allow them to come from an outside source not part of the transformer—they may thus carry information not computable by the model; see §˜D.1 for additional discussion.
@@ 5 Conclusion
Appendix A Related work on the experssivity of MDMs
@@
Feng et al. (2025) analyze the implications of the factorized backward process on the expressivity of MDMs. They find that MDMs can approximate n-gram LM distributions arbitrarily well with constantly many sampling steps and linearly many sampling steps suffice to approximately generate from any regular language. They also provide a linear lower bound on the number of sampling steps required to capture the support of general regular languages—a consequence of the assumed factorization, which is incompatible with the sequential generation of regular languages. Crucially, this differs from our recognition setting in which the string to be processed is given and only its membership decision has to be generated.
@@ B.1 Notation
We denote the concatentation of two strings ${{\bm{w}}}_{1},{{\bm{w}}}_{2}\in{{{{{{\Sigma}}^{*}}}}}$ as ${{\bm{w}}}_{1}\circ{{\bm{w}}}_{2}$ or simply ${{\bm{w}}}_{1}{{\bm{w}}}_{2}$. We define the intereleaving of the vectors ${{{\bm{x}}}},{{{\bm{y}}}}\in{{\mathbb{R}}}^{{D}}$ as ${{{{{\bm{x}}}}}^{\frown}{{{{\bm{y}}}}}}\in{{\mathbb{R}}}^{2{{D}}}$ where
@@ B.4.4 Sampling from a (transformer) language model
Let ${{{\bm{l}}}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{{{{\bm{E}}}}}}{{{{{\bm{h}}}}}}$ for ${{{{{\bm{h}}}}}}\in{{\mathbb{R}}}^{{{D}}}$ and ${{{{{\bm{E}}}}}}\in{{\mathbb{R}}}^{{{|{{\overline{{{\Sigma}}}}}|}}\times{{D}}}$ be the logits of a probability distribution over $\overline{{{\Sigma}}}$}}. The Gumbel-max sampling from is defined as
@@ B.5 Chain-of-thought reasoning
To facilitate a more convenient connection to MDMs, we introduce a parallel chain-of-thought process that can be seen as a generalization of CoT transformers that generates multiple symbols at once.
@@ B.6 Padded looped transformers
The dynamic computational depth of PLTs endows them with the ability to perform more complex reasoning tasks by iteratively refining their hidden states over multiple timesteps. Importantly, these reasoning steps include both sequential and parallel processing of the input symbols, allowing for both parallel efficiency as well as depth in the reasoning process.
@@ C.1 The planner as a principled departure from uniform unmasking
The choice to allow the MDM to resample already unmasked symbols also departs from the analytical reverse process of MDMs, which only unmask masked symbols. This is, however, a principled choice that allows the MDM to correct earlier mistakes and is supported by empirical evidence that resampling already unmasked symbols can improve generation quality (von Rütte et al., 2025). Practically, this choice enables conciser connections to PLTs and pCoT transformers. Moreover, all results in this work can easily be adapted to MDMs with mask dominated planners that do not resample already unmasked symbols. This is justified by the following theorem that shows that any MDM with an unrestricted planner can be simulated by an MDM with a mask dominated planner by increasing the output space by a factor of T{T} to account for the inability to resample. In the following, we refer to MDMs with mask dominated planners as simple MDMs (sMDMs).

### v2

@@ Abstract
Masked diffusion models (MDMs) for text offer a compelling alternative to traditional autoregressive language models. Parallel generation makes them efficient, but their computational capabilities and the limitations inherent in their parallelism remain largely unexplored. To this end, we characterize what types of reasoning problems MDMs can provably solve and how efficiently. We do this by connecting MDMs to the well-understood reasoning frameworks of chain of thought (CoT) and padded looped transformers (PLTs) in the finite-precision log-width setting: We show that MDMs and polynomially-padded PLTs are, in fact, equivalent in this setting, and that MDMs can solve all problems that CoT-augmented transformers can. Moreover, we showcase classes of problems (including regular languages) for which MDMs are inherently more efficient than CoT transformers, where parallel generation allows for substantially faster reasoning.
@@ 1 Introduction
Many complex problems can be decomposed into smaller, independent sub-problems, making them naturally suited for parallel computation. For example, we can compute the value of a mathematical expression by evaluating its sub-expressions independently and combining the results (see Fig.˜3). However, dominant autoregressive language models (LMs) tackle these problems sequentially. Methods like chain of thought (CoT), for instance, generate solutions one step at a time, failing to capitalize on the underlying parallel structure. Parallel generation by masked diffusion models (MDMs) offers a compelling alternative. Recent advances have positioned MDMs as a viable contender to autoregressive LMs in language modeling, code generation, and even molecule design (Lou et al., 2024; Zhang et al., 2025; Sun et al., 2025). However, the fundamental reasoning capabilities of MDMs remain poorly understood, which limits the extent to which we can leverage their potential and apply them to appropriate tasks. This work bridges that gap by providing the first formal characterization of the expressivity of MDMs, clarifying their fundamental computational strengths and weaknesses.
Masked Diffusion Models Padded Looped Transformers Chain of Thought Added stochasticityThm.˜3.1Extra output spaceThm.˜3.1Extra stepsThm.˜3.4Extra output spaceThm.˜3.3Sequentiality Bottleneck under $\log{{N}}$ steps.Cor.˜3.7
Figure 1: A summary of masked diffusion model expressivity in relation to padded looped transformers and chain of thought. ${{{\mathcal{X}}}}\hookrightarrow{{{\mathcal{Y}}}}$ indicates the inclusion of $\mathcal{X}$}}} in $\mathcal{Y}$}}}. Dashed arrows represent strict inclusions. Red arrows denote novel results. Edge labels with blue background indicate constraints on the source node, while labels with orange background indicate constraints on the target node.
@@
Figure 2: A summary of masked diffusion model expressivity in relation to classical complexity classes. ${{{\mathcal{X}}}}\hookrightarrow{{{\mathcal{Y}}}}$ indicates the inclusion of $\mathcal{X}$}}} in $\mathcal{Y}$}}}, and ${{{\mathcal{X}}}}\leftrightarrow{{{\mathcal{Y}}}}$ indicates equality. Dashed arrows represent strict inclusions. Red arrows denote novel results. $\mathtt{Reg}$ refers to all regular languages. Edge labels with blue background indicate constraints on the source node, while labels with orange background indicate constraints on the target node (in case of $\leftrightarrow$, on MDMs).
@@
Our analysis builds upon the study of LM expressivity—the formal characterization of the problems whose solution the neural architecture of an LM (with appropriate parameters) can express. While this field comprehensively describes autoregressive LMs, its findings do not directly apply to MDMs due to their fundamentally different, non-sequential processing of text, leaving the theoretical studies of these two paradigms largely disconnected. Prior theoretical work on MDMs has focused on the limitations of their factorized backward process and its convergence properties. This research has shown that while MDMs can approximate $n$-gram LMs with constantly many denoising steps, the number of steps must grow linearly with the string length even for simple LMs such as probabilistic regular languages (Feng et al., 2025; Li & Cai, 2025). [^1] We provide a detailed overview of related work in App. A. However, the asymptotic nature of these findings, together with the strict assumptions on the theoretical model, makes it difficult to draw concrete conclusions about the practical reasoning capabilities of MDMs, leaving a critical gap: A theoretical framework for MDMs that is both (1) formally rigorous enough to provide a comprehensive picture of how MDMs can use the combination of parallelism and iterative refinement for formal reasoning, and (2) faithful to how they are implemented in practice. Our work introduces such a framework, providing a tight characterization of their reasoning capabilities.
@@
We refer to the fact that CoT cannot take advantage of this parallelism as the sequentiality bottleneck of CoT. Takeaway˜2 highlights the potential efficiency gains of parallel CoT that generates multiple symbols at once. Our analysis, in fact, identifies a tighter and more natural connection between MDMs and this variant, which we refer to as pCoT. The parallelism of MDMs also facilitates a close connection with looped and padded transformers, where looping naturally maps to denoising steps and padding tokens to the generated tokens of an MDM. We find that the class of problems solvable by MDMs is, in fact, precisely equivalent to that solvable by padded looped transformers.
@@
We summarize these takeaways and their relationships in Figs.˜1 and 2. Our proofs emphasize the affordances and difficulties in solving reasoning problems with MDMs. We find, for example, that the discrete nature of the generated text, which serves as a communication channel between individual denoising steps, limits the amount of information that can be passed between steps. This necessitates extra output space to store the intermediate computations and is analogous to the discrepancy between fixed-precision and log-precision transformers (Li et al., 2024), and the difference between the classes $\mathtt{AC}$}} and $\mathtt{TC}$}}, the class of circuits with threshold gates. We find that positional encodings that carry information not computable by transformers themselves are crucial in locating this information. We also observe that, while unmasked attention makes MDM attention patterns flexible, it complicates left-to-right processing, which is often natural in human language—we find that exact implementation of causal masking requires quadratically more output space.
@@ 2.2.1 Our idealization of masked diffusion models
We aim to understand the expressivity of the reverse process. To this end, we introduce an idealization that captures its key aspects—iterative unmasking and infilling—and provides a principled lens for understanding the expressivity of practical MDMs and comparing them to well-known paradigms such as CoT. Here, we describe the high-level ideas; see §˜B.7 for the full formal model.
@@
We denote the class of our idealized MDMs by $\mathtt{MDM}$}. For a more detailed discussion, including the theoretical connection of our idealization to existing MDM variants, see App.˜C.
@@ 3.2 MDMs and CoT can (inefficiently) simulate each other
The simulation incurs a quadratic blow-up in the padding length. This is not due to an inherent feature of the diffusion process but rather the challenge of simulating masked attention with unmasked one (cf. Lem.˜D.14). [^13] To the best of our knowledge, Lem. D.14 is the first result showing how to simulate masked attention with unmasked attention and might be of interest in its own right. In particular, if the MDM transformer is causally masked, the blow-up disappears. Moreover, if the unmasked transformer can simulate masking more efficiently (with, for example, more expressive scoring functions), the blow-up can be alleviated. [^14] This is, for example, used by Saunshi et al. (2025, Thm 5.4) to show that unmasked PLTs can simulate CoT with no blow-up in the padding using a masking function—an additional step in the computation of attention scores that allows for zeroing out of irrelevant keys—and with linearly-increasing width. While Lem.˜D.14 could possibly be improved, it is interesting to note that the seemingly more general, unmasked nature of MDMs might negatively impact their ability to align with human-oriented sequential processing captured by causal masking, which could provide a useful inductive bias for the masked models. [^15] Interestingly, some existing work finds that MDMs that decode based on the most confident symbols tend to decode autoregressively (Gong et al., 2025). In this sense, the unmasked nature of MDMs could be seen as a hurdle that the model has to overcome to eventually rely on more autoregressive generation. This further motivates the development of hybrid models that combine autoregressive generation of entire blocks with non-autoregressive infilling within the blocks (Nie et al., 2025; Arriola et al., 2025; Song et al., 2025, inter alia).
@@ 4 Discussion
In other words, these problems, in general, require a “CoT-style” step-by-step sequential solution. Due to the overhead introduced by unmasked attention of MDMs (such as the inability to maintain KV-cache), such problems are more efficiently solved by standard autoregressive CoT transformers.
@@
We strive to compare different paradigms fairly by analyzing a specific implementation of MDMs—one based on a specific idealization of transformers. Some inherent differences between the frameworks, however, remain. One is the dichotomy between using causal masking for CoT and unmasked transformers for MDMs and PLTs. We focus on unmasked models to analyze the more natural and popular implementations rather than artificially constraining MDMs and PLTs to causal masking.
Positional encodings.
Another impactful decision is the nature of positional encodings (PEs). While we assume relatively simple PEs standard in theoretical literature (in particular, logspace-computable, cf. §˜B.4), we allow them to come from an outside source not part of the transformer—they may thus carry information not computable by the model. This highlights a crucial feature: The computational capability in the ${\mathtt{L}}\textnormal{-uniform}$} transformer family is deliberately split between the PEs and the transformer’s computation mechanism. Fixed-precision transformers are limited to a subset of ${{\mathtt{AC}}}^{\mathtt{0}}$}}}} and cannot compute division, modulo, or binary representations (Li et al., 2024). The PEs provide this missing information: Binary encodings of position/length and results of division/modulo operations (see §˜D.1). The transformer mechanism can then perform ${{\mathtt{AC}}}^{\mathtt{0}}$}}}} operations (addition, thresholding, multiplication by powers of two) on this pre-computed data. Thus, in Thm.˜3.2, the transformer is looking up pre-computed (but still simple, log-space) information from the PEs rather than computing it. When chaining decoding steps, modulo and division information is needed for correct indexing—making ${\mathtt{L}}\textnormal{-uniform}$} PEs a prerequisite for “looping” to increase computational power beyond fixed-depth models. We emphasize that offloading computation to PEs is standard in the expressivity literature (Li et al., 2024; Saunshi et al., 2025; London & Kanade, 2025), often encoding entire circuits to be simulated. In contrast, our PEs are simpler: Functions of input position and length, not circuit-specific.
Relaxing assumptions.
Our constructions rely on a planner that perfectly unmasks the input string and an unmasker that models the conditional distributions perfectly. These are idealizations of confidence-based planners and learned unmaskers. How a trained planner can learn to approximate this ideal behavior is an important question—one of learnability rather than pure expressivity. As is standard in expressivity analysis of language models, expressivity is a necessary first step before one can analyze how a model can learn to represent such functions. While learnability of formal languages is an active area of research, we are not aware of any existing results that would characterize the learnability of functions required for planning (e.g., modulo and division) apart from the issue of the sensitivity of modular counting (Hahn & Rofin, 2024), which suggests that learning planning behaviour used in our constructions is difficult for transformers.
Empirical validation.
The main purpose of this work is largely theoretical. It provides a framework for understanding the expressivity of MDMs in relation to other popular paradigms and provides new insights into the tasks that MDMs are well-suited for. In fact, this analysis is motivated by existing empirical results that show MDMs outperforming autoregressive models on tasks that allow for parallel reasoning (Kim et al., 2025; Ye et al., 2025). The lack of literature on training MDMs on formal languages, however, makes it difficult to draw concrete empirical implications of the presented results. Thus, we see designing targeted experiments and benchmarks that empirically explore the theoretical benefits of MDMs as an important direction for future work.
@@ 5 Conclusion
Acknowledgments
Anej Svete is supported by the ETH AI Center doctoral fellowship.
Appendix A Related work on the expressivity of MDMs
@@
Feng et al. (2025) analyze the implications of the factorized backward process on the expressivity of MDMs. They find that MDMs can approximate n-gram LM distributions arbitrarily well with constantly many sampling steps and linearly many sampling steps suffice to approximately sample from any regular language. They also provide a linear lower bound on the number of sampling steps required to capture the support of general regular languages—a consequence of the assumed factorization, which is incompatible with the sequential generation of regular languages. Crucially, this differs from our recognition setting in which the string to be processed is given and only its membership decision has to be generated.
@@ B.1 Notation
We denote the concatenation of two strings ${{\bm{w}}}_{1},{{\bm{w}}}_{2}\in{{{{{{\Sigma}}^{*}}}}}$ as ${{\bm{w}}}_{1}\circ{{\bm{w}}}_{2}$ or simply ${{\bm{w}}}_{1}{{\bm{w}}}_{2}$. We define the interleaving of the vectors ${{{\bm{x}}}},{{{\bm{y}}}}\in{{\mathbb{R}}}^{{D}}$ as ${{{{{\bm{x}}}}}^{\frown}{{{{\bm{y}}}}}}\in{{\mathbb{R}}}^{2{{D}}}$ where
@@ B.4.4 Sampling from a (transformer) language model
Let ${{{\bm{l}}}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{{{{\bm{E}}}}}}{{{{{\bm{h}}}}}}$ for ${{{{{\bm{h}}}}}}\in{{\mathbb{R}}}^{{{D}}}$ and ${{{{{\bm{E}}}}}}\in{{\mathbb{R}}}^{{{|{{\overline{{{\Sigma}}}}}|}}\times{{D}}}$ be the logits of a probability distribution over $\overline{{{\Sigma}}}$}}. The Gumbel-max sampling is defined as
@@ B.5 Chain-of-thought reasoning
To facilitate a more convenient connection to MDMs, we introduce a parallel chain-of-thought process that can be seen as a generalization of CoT transformers, one that generates multiple symbols at once.
@@ B.6 Padded looped transformers
The dynamic computational depth of PLTs endows them with the ability to perform more complex reasoning tasks by iteratively refining their hidden states over multiple timesteps. Importantly, these reasoning steps include both sequential and parallel processing of the input symbols, allowing for both parallel efficiency and depth in the reasoning process.
@@ C.1 The planner as a principled departure from uniform unmasking
The choice to allow the MDM to resample already unmasked symbols also departs from the analytical reverse process of MDMs, which only unmask masked symbols. This is, however, a principled choice that allows the MDM to correct earlier mistakes and is supported by empirical evidence that resampling already unmasked symbols can improve generation quality (von Rütte et al., 2025). In practice, this choice enables more concise connections to PLTs and pCoT transformers. Moreover, all results in this work can easily be adapted to MDMs with mask dominated planners that do not resample already unmasked symbols. This is justified by the following theorem that shows that any MDM with an unrestricted planner can be simulated by an MDM with a mask dominated planner by increasing the output space by a factor of T{T} to account for the inability to resample. In the following, we refer to MDMs with mask dominated planners as simple MDMs (sMDMs).